# A General Model for Static Contact Angles

### v1

@@ 3 Results and Discussion
Figure 4: Micro/nanocavity arrays. Local static contact angle $\theta$  predicted by Eq. 7 as a function of liquid-vapor surface fractions $\phi_{B}$ (bulk) and $\phi_{A}$ (ambient) adjacent to the contact line, for three Young angles: $\theta_{Y}=40^{\circ},\,90^{\circ},\,\&\,120^{\circ}$. The results correspond to surfaces with sharp topographic features, with a small solid-top area fraction $\varphi_{S}=0.8$ and for two different lateral-plus-basal area fractions $\Lambda=1$ (top row) and $\Lambda=2$ (bottom row). Wetting states: (W) Wenzel, (CB) Cassie-Baxter, (IW) Inverse Wenzel, and (H) Hemiwicked.
@@ 4 Conclusions
This work presents a unified and simply to use model for predicting static contact angles that extends Young’s local thermodynamic framework to surfaces with general topography and non-uniform local liquid infiltration. The formulation makes explicit the geometric controls through the solid-top area fraction $\varphi_{S}$ and the normalized lateral plus basal area $\Lambda$, and it resolves the local wetting state adjacent to the contact line with two side specific infiltration fractions $(\phi_{A},\phi_{B})$. In closed form, the model recovers the classical limits of Wenzel, Cassie-Baxter, and hemiwicking, and it predicts a previously unreported limiting configuration with practical relevance, an Inverse Wenzel state characterized by a Cassie core coexisting with a hemiwicked film. Because the inputs are measurable from surface geometry and standard force tensiometer data, the framework links morphology, local infiltration, and apparent angle, and it provides sharp bounds for advancing and receding states.
@@
Analytical exploration across representative surface morphologies yields design guidelines. For sharp random relief with very small $\varphi_{S}$ and large $\Lambda$, the model predicts broad tunability: robust superhydrophilicity with $\theta\to 0$ on intrinsically wettable chemistries when the bulk side is liquid filled ($\phi_{B}=0$), with weak dependence on the ambient side condition. The same class of morphologies can also support very large apparent angles on hydrophilic substrates through the Inverse Wenzel state if the liquid is nearly fully suspended over trapped vapor on the bulk side ($\phi_{B}\gtrsim 0.9$), which highlights how large $\Lambda$ amplifies $\theta$. For engineered pillar arrays with moderate $\varphi_{S}\sim 0.5$, sensitivity to $(\phi_{A},\phi_{B})$ is reduced, so achieving superhydrophilicity requires taller features and low $\theta_{Y}$, while superhydrophobicity appears only in a narrow window of side specific infiltration. For cavity rich surfaces with large $\varphi_{S}\sim 0.8$, superhydrophilicity is attainable for hydrophilic chemistries when $\Lambda>1$ and the ambient side wets appreciably, whereas superhydrophobic states are limited to intrinsically hydrophobic chemistries and restricted ranges of $(\phi_{A},\phi_{B})$.
The model also strengthens connections to near equilibrium (quasi-static) dynamics. By predicting feasible ranges of apparent angles and hysteresis for given morphology, chemistry, and ambient conditions, it provides a basis to rationalize the slow quasi static spreading of droplets and capillary imbibition reported for glass and related materials 58; 16; 51; 59; 52; 53; 18. The side specific infiltration parameters $(\phi_{A},\phi_{B})$ capture metastable pinning and depinning pathways near the contact line, thus informing the control of entry pressures, relaxation times, and near-equilibrium flow thresholds. The same predictions constrain the capillary energy landscape that governs the quasi static relaxation of micro and nanoparticles at liquid fluid and liquid solid interfaces 20; 60; 61; 62; 63; 24; 64; 65, enabling surface design strategies, for example increasing $\Lambda$ at fixed $\varphi_{S}$, to accelerate or arrest adsorption and to tune dissipation during interfacial relaxation.
Future experimental works can use force versus displacement measurements with the Wilhelmy plate method to calibrate, test, and extend the model under controlled surface chemistry, cleanliness, ambient humidity, and temperature. The direct relation between the measured force and the apparent angle along advancing and receding branches enables inference of the side specific infiltration fractions $(\phi_{A},\phi_{B})$, validation of the predicted bounds and transitions, and targeted searches for the proposed Inverse Wenzel state. In parallel, the framework can be closed with thermodynamic relations that predict $\phi_{B}$ and $\phi_{A}$ from measurable quantities such as Laplace pressure, interfacial work, and local free energy differences, for example by minimizing the interfacial free energy of a representative unit cell at the prescribed reference height given $\theta_{Y}$, $\varphi_{S}$, and $\Lambda$. The combination of Wilhelmy data and these closures will enable parameter free predictions of static angles and hysteresis and guide refinement of the model in regimes where kinetic barriers and condensation effects become important.

### v2

@@ 2.3 Micro/nanocavity arrays
Figure 4: Micro/nanocavity arrays. Local static contact angle $\theta$  predicted by Eq. 7 as a function of vapor fractions $\phi_{B}$ (bulk) and $\phi_{A}$ (ambient) adjacent to the contact line, for three Young angles: $\theta_{Y}=40^{\circ},\,90^{\circ},\,\&\,120^{\circ}$. The results correspond to surfaces with sharp topographic features, with a small solid-top flat area fraction $\varphi_{S}=0.8$ and for two different lateral-basal area fractions $\Lambda=1$ (top row) and $\Lambda=2$ (bottom row). Wetting states: (W) Wenzel, (CB) Cassie-Baxter, (IW) Inverse Wenzel, and (H) Hemiwicked.
The final cases analyzed (cf. Fig. 4) correspond to a large top-solid flat area fraction, $\varphi_{S}=0.8$, and lateral-basal area fractions $\Lambda=1$ and $\Lambda=2$. This type of surface topography can be attained, for example, by patterning a smooth and flat surface with arrays of cavities with controllable micro- or nanoscale dimensions (e.g., Refs. 55; 56; 57). Note that if the cavities are not interconnected, capillary condensation and/or direct prewetting of the surface, rather than hemiwicking, are feasible mechanisms for infiltrating the surface topography. For this type of topography, robust superhydrophilicity with $\theta\to 0$ can be realized for $\phi_{A}\gtrsim 0.5$ and $\phi_{B}\lesssim 0.5$ in the case of large lateral-basal areas with $\Lambda>1$ and hydrophilic substrates with $\theta_{Y}\lesssim 40^{\circ}$. Superhydrophobicity is more difficult to achieve with this type of surface topography, and $\theta\geq 150^{\circ}$ is predicted only for highly hydrophobic substrates when $\phi_{A}\gtrsim 0.8$ and $\phi_{B}\lesssim 0.5$, which represent a rather narrow range of conditions.
@@ 3 Conclusions
This work introduces a unified model for static contact angles that extends Young’s local thermodynamic framework to surfaces with micro- or nanoscale topographic features and nonuniform liquid infiltration near the contact line. The model employs as input parameters the solid-top flat area fraction $\varphi_{S}$ and the lateral plus basal area $\Lambda$, and considers side-specific liquid infiltration under the liquid bulk and ambient phase through the vapor fractions $\phi_{A}\in[0,1]$ and $\phi_{B}\in[0,1]$. The unified model recovers the Wenzel, Cassie-Baxter, and hemiwicking limits, and it predicts a fourth limiting configuration of practical interest, an Inverse Wenzel state with a Cassie-like core and a uniform liquid film under the ambient phase. Because all inputs can be obtained from surface measurements and standard force tensiometer data, the model can link geometric parameters, liquid infiltration degree, and static contact angles, and it yields clear bounds for advancing and receding states and thus for static hysteresis.
@@
Analytical predictions from the unified model provide direct design guidance for contact angle control under varying interfacial and environmental conditions. For sharp random surface features with very small top-solid flat area fractions $\varphi_{S}<0.1$ and large lateral-basal areas $\Lambda>1$, the model predicts robust superhydrophilicity with $\theta\to 0$ on intrinsically wettable substrates ($\theta_{Y}<90^{\circ}$) when the bulk side is liquid filled ($\phi_{B}=0$), for arbitrary ambient-side conditions ($0\leq\phi_{A}\leq 1$). Notably, the same class of sharp topographic features can also support large contact angles $\theta\gtrsim 150^{\circ}$ on hydrophilic substrates through the Inverse Wenzel state if the liquid is nearly fully suspended over kinetically or physically trapped vapor on the bulk side ($\phi_{B}\gtrsim 0.9$), which highlights how large lateral areas $\Lambda>1$ can amplify variations of the attainable static contact angle $\theta$ and produce extremely large hysteresis ranges. For engineered pillar arrays with moderate $\varphi_{S}\sim 0.5$, sensitivity to local liquid/vapor infiltration $(\phi_{A},\phi_{B})$ and thus ambient conditions is reduced, and achieving superhydrophilicity requires tall features ($\Lambda>1$) and low Young contact angles, while superhydrophobicity appears only for very large Young contact angles $\theta_{Y}\geq 120^{\circ}$ (e.g., for silanized substrates) within a narrow window of side-specific vapor fractions $(\phi_{A}\gtrsim 0.9,\phi_{B}<0.5)$. For surfaces with large $\varphi_{S}\sim 0.8$, such as those produced by micro/nanoscale cavity arrays, superhydrophilicity is robustly attainable for hydrophilic substrates when $\Lambda\geq 1$ and the ambient side is partially infiltrated by liquid, whereas attaining high contact angles is limited to intrinsically hydrophobic surfaces and moderately narrow range of vapor/liquid infiltration fractions $(\phi_{A}\gtrsim 0.8,\phi_{B}\lesssim 0.5)$.
The model predictions also help account for the slow near-equilibrium relaxation that arises from transitions between numerous metastable states when a contact line moves over surfaces with random nanoscale topographies. In droplet spreading, capillary imbibition, and nanoparticle adsorption, relaxation proceeds through a sequence of quasi-static configurations where the contact line remains locally pinned while the local static contact angle can vary over a wide range of values. It is therefore possible that both the geometry of topographic features and environmental conditions controlling local variations in $\phi_{A}$ and $\phi_{B}$ prescribe the relaxation rates documented in prior works. 19; 20; 58; 16; 59; 51; 52; 53; 18
The theoretical framework presented in this work suggests concrete directions for further theoretical developments and experimental verification. Future developments of the presented unified model may include a thermodynamic closure for $\phi_{A}$ and $\phi_{B}$, by considering the Laplace, Kelvin, and disjoining pressure, along with capillary entry thresholds determined from topographic analysis (e.g., AFM, SEM, or profilometry); thereby making the vapor fractions $(\phi_{A},\phi_{B})$ predictable state variables rather than free model parameters. Additionally, the unified model can be readily applied for complex surface topographies, with overhangs or hierarchical features, for which $\varphi_{L}+\varphi_{S}>1$. In such cases, $\varphi_{L}$ should be obtained from the level set area at an appropriate reference height or from three-dimensional surface reconstructions. The unified model in Eq. 7 is generally valid when the properly determined values of $\varphi_{L}$ and $\Lambda$ are used. Force-displacement measurements with the Wilhelmy plate method at low displacement rates can be used to determine advancing and receding contact angles and infer $(\phi_{A},\phi_{B})$ along the contact line path. This type of experimental analysis could validate and refine analytical predictions for the static contact angle and hysteresis range under different environmental conditions, identify transitions between limiting and intermediate states, and guide the potential realization of the Inverse Wenzel configuration.